Miandad scored 85 runs in the 17th innings and thereby increased the run average of the previous 16 innings by 3.
What were his average runs in 17 innings?

Answer by stanbon(75887)

Miandad scored 85 runs in the 17th innings and thereby increased the run average of the previous 16 innings by 3.
What were his average runs in 17 innings?
---------------
Let 16 innings average be "x".
Then number of runs scored in the 16 innings is 16x
----------
(16x + 85)/17 = x + 3
16x + 85 = 17x + 51
----
x = 34
---
Ans: 17 innings average = 37
--------------------
Cheers,
Stan H.
-------------